# DecAlign: Hierarchical Cross-Modal Alignment for Decoupled Multimodal Representation Learning

### v1

@@ Abstract
Multimodal representation learning aims to capture both shared and complementary semantic information across multiple modalities. However, the intrinsic heterogeneity of diverse modalities presents substantial challenges to achieve effective cross-modal collaboration and integration. To address this, we introduce DecAlign, a novel hierarchical cross-modal alignment framework designed to decouple multimodal representations into modality-unique (heterogeneous) and modality-common (homogeneous) features. For handling heterogeneity, we employ a prototype-guided optimal transport alignment strategy leveraging gaussian mixture modeling and multi-marginal transport plans, thus mitigating distribution discrepancies while preserving modality-unique characteristics. To reinforce homogeneity, we ensure semantic consistency across modalities by aligning latent distribution matching with Maximum Mean Discrepancy regularization. Furthermore, we incorporate a multimodal transformer to enhance high-level semantic feature fusion, thereby further reducing cross-modal inconsistencies. Our extensive experiments on four widely used multimodal benchmarks demonstrate that DecAlign consistently outperforms existing state-of-the-art methods across five metrics. These results highlight the efficacy of DecAlign in enhancing superior cross-modal alignment and semantic consistency while preserving modality-unique features, marking a significant advancement in multimodal representation learning scenarios. Our project page is at https://taco-group.github.io/DecAlign/ and the code is available at https://github.com/taco-group/DecAlign.
@@ 1 Introduction
Multimodal representation learning seeks to effectively harmonize diverse modalities by capturing their shared semantics while retaining modality-unique characteristics. This approach has achieved significant advances across various domains such as multimodal sentiment analysis [18, 4, 37], recommendation systems [24, 23], autonomous driving [13, 44, 26, 43], out-of-distribution detection [5, 16], and large vision language models [45, 38, 56]. Despite these successes, the inherent modality heterogeneity—arising from distinct data distributions, representation scales, and semantic granularities—remains a significant barrier that limits their potential for seamless cross-modal collaboration.
Figure 1: Overview of the proposed DecAlign framework, where multimodal representations capture both modality-unique characteristics and shared commonalities. By decoupling them into modality heterogeneity and homogeneity, followed by a hierarchical alignment strategy, we achieve seamless alignment from heterogeneous to homogeneous features.
This challenge is further intensified by the entanglement of modality-unique (heterogeneous) patterns and cross-modal common (homogeneous) semantics. Conventional multimodal fusion methods typically simplify the issue by projecting raw multimodal data into unified spaces via straightforward concatenation or linear transformations [10, 54]. However, this indiscriminate fusion entangles modality-unique features with shared semantics, leading to semantic interference, where detailed unimodal characteristics may disrupt global cross-modal relationships [19, 46]. This phenomenon is particularly evident when dealing with dimensional mismatches—such as high-dimensional, spatially correlated image features paired with low-dimensional, temporally correlated text features [41, 40, 58]. Such mismatches hinder effective cross-modal semantic alignment, causing potential information redundancy or loss during fusion.
@@
Figure 2: DecAlign achieves superior performance compared to state-of-the-art methods across multiple multimodal benchmarks. The bubble size represents relative model performance, illustrating the trade-off between Acc-2 and F1 Score.
To overcome these limitations, we propose DecAlign, a hierarchical cross-modal alignment framework for multimodal representation learning. As illustrated in Figure 1, DecAlign first explicitly decouple multimodal representations into modality-unique (heterogeneous) and modality-common (homogeneous) features through specialized encoders. Then, leveraging a dual-stream cross-modal alignment mechanism, DecAlign individually handles modality characteristics at different granularities: ❶ For heterogeneity, we propose prototype-based optimal transport alignment [28] using Gaussian Mixture Modeling (GMM) [2] and multi-marginal transport plans [27], effectively mitigating distribution discrepancies and constrain modality-unique interference. Furthermore, we incorporate a multimodal transformer module with cross-modal attention to align high-level semantics, bridges distribution gaps, reinforces discriminative representations, bridging semantic gaps across modalities. ❷ For homogeneity, DecAlign achieves semantic consistency via latent distribution matching and Maximum Mean Discrepancy (MMD) regularization. Finally, we concatenate the transformer-aligned modality-unique features with modality-common features, passing them through a learnable fully connected layer for downstream tasks. The key contributions of ourwork are summarized as follows:
@@
  • Effectiveness. Through comprehensive evaluation on four benchmark datasets, we demonstrate that DecAlign consistently surpasses state-of-the-art methods, confirming its efficacy for multimodal representation learning.
Figure 3: The framework of our proposed DecAlign approach, illustrated in a multimodal setting with visual, audio, and language inputs. Modality Feature Encoders first extract unimodal embeddings, which are then decoupled into modality heterogeneous and homogeneous components by modality-unique/common encoders. Heterogeneous features are aligned via optimal transport-based cross-modal prototypes, and homogeneous semantics are aligned through latent space semantics and Maximum Mean Discrepancy-based distribution matching. Heterogeneous features are refined by multimodal transformer for capturing finer-grained cross-modal interactions, then concatenated with homogeneous features and passed through a fully connected layer for downstream tasks.
2 Related Work
2.1 Multimodal Representation Learning
Multimodal representation learning aims to integrate heterogeneous data from diverse modalities into a cohesive framework that captures complementary semantic information [29, 20, 1]. Recent methods have achieved significantly performance improvements by leveraging representation-based and cross-modal interaction approaches. Specifically, Self-MM [50] applies self-supervised contrastive learning and masked modeling to enhance the mutual information across modalities, HGraph-CL [22] introduces hierarchical graph contrastive learning to model intricate interactions across modalities. However, the heterogeneity and complementary information inherent in multimodal representations are intrinsically entangled, making it challenging to fully harness their complementary strengths while preserving their unique characteristics. Inspired by this insight, MISA [11] separates multimodal representations into modality-invariant and modality-unique features with contrastive and reconstruction losses, DMD [17] further introduces graph cross-modal knowledge distillation to explicitly model the correlations across modalities. However, existing methods are often constrained to modeling modalities from a global perspective, overlooking the token-level local semantic inconsistencies that arise in cross-modal interactions. Our proposed DecAlign enables fine-grained multimodal representation learning through hierarchical alignment, progressing from local to global and heterogeneity to homogeneity, ensuring precise cross-modal integration and semantic consistency.
2.2 Cross-Modal Alignment
@@
The core challenge in multimodal tasks lies in the inherent heterogeneity across modalities [58], characterized by structural, distributional, and semantic disparities, which restricts the effective synergy of multimodal homogeneous features. To address this, existing solutions can be broadly categorized as follows: ❶ Shared Representation, which aims to learn a unified latent space for cross-modal semantic consistency. For example, CLIP-based methods [30, 8] use contrastive learning to align image-text pairs in a shared embedding space, while Uni-Code [42] employs cross-modal information disentangling and exponential moving average to align semantically equivalent information in a shared latent space. ❷ Transformer-based methods that apply cross-attention to dynamically capture key information in cross-modal interactions [35, 47, 12]. ❸ Modality Translation, which establishes mappings between modalities through cross-modal generation or reconstruction [25, 53, 33]. ❹ Cross-Modal Knowledge Distillation, which addresses inter-modal contribution imbalances and explores cross-modal correlations. For example, DMD [17] employs graph distillation for dynamic knowledge transfer, and UMDF [15] uses unified self-distillation to learn robust representations from consistent multimodal distributions. Unlike methods that overemphasize homogeneous information, we tackle the issue of over-alignment diminishing modality-unique characteristics through representation decoupling and a hierarchical alignment mechanism, ensuring cross-modal semantic consistency while retaining unimodal characteristics.
3 The Proposed Method
@@ 3.1 Multimodal Feature Decoupling
Given a multimodal dataset with $M$ modalities, we first employ 1D temporal convolutional layers on unimodal features extracted from modality-unique encoders to aggregate temporal information. This shallow encoding process preserves the original temporal dimension while unifying the feature dimensions across modalities. The resulting unimodal features are expressed as: $\tilde{\mathbf{X}}_{m}\in\mathbb{R}^{T_{s}\times d_{s}}$, where $m$ denotes the modality index, $T_{m}$ represents the temporal dimension, and $d_{m}$ is the feature dimension of the corresponding modality.
The primary challenge in multimodal tasks lies in the inherent heterogeneity across modalities, hindering the integration of homogeneous features. To address this, we decouple the multimodal representations into modality-common features, which emphasize semantic consistency across modalities, and modality-unique features, capturing modality-unique characteristics with some redundancy. Building upon this, we employ three modality-unique encoders $\mathbf{E}_{\mathrm{uni}}^{(m)}$ and a modality-shared encoder $\mathbf{E}_{\mathrm{com}}$, to extract heterogeneous features as $\mathcal{F}_{\mathrm{uni}}^{(m)}=\mathbf{E}_{\mathrm{uni}}^{(m)}(\tilde{\mathbf{X}}_{m})$ and cross-modal homogeneous features as $\mathcal{F}_{\mathrm{com}}^{(m)}=\mathbf{E}_{\mathrm{com}}(\tilde{\mathbf{X}}_{m})$.
Considering the inherent heterogeneity and potential redundancy across modalities, we refine the decoupling process by clearly separating modality-unique and modality-common features. Instead of explicitly modeling distributions or computing mutual information which can significantly increase computational complexity, we use cosine similarity to quantify their potential overlap. Hence, the loss of decoupling process is formally defined as:
$\mathcal{L}_{dec}=\sum_{m=1}^{M}\frac{\mathcal{F}_{\mathrm{uni}}^{(m)}\cdot(\mathcal{F}_{\mathrm{com}}^{(m)})^{\mathrm{T}}}{||\mathcal{F}_{\mathrm{uni}}^{(m)}||~{}||\mathcal{F}_{\mathrm{com}}^{(m)}||}$ (1)
@@ 3.2 Heterogeneity Alignment
In multimodal tasks, modality-unique features encapsulate distinct characteristics inherent to each modality. However, differences in representation forms and distributions across modalities can result in semantic shifts during direct alignment, ultimately weakening the robustness of cross-modal alignment. To address this, we refine local alignment using category prototypes and leverage multi-marginal optimal transport-based dynamic transfer mechanism for adaptive, fine-grained feature alignment across modalities.
3.2.1 Prototype Generation
To flexibly capture the complex distributions and potential correlations in multimodal data, we employ the Gaussian Mixture Model (GMM), which leverages its soft assignment mechanism and Gaussian distribution assumption to more accurately represent the prototype structures of different modality features. We first model modality-unique features using GMM, with prototypes represented by the mean and covariance of Gaussian distributions:
@@ 3.2.1 Prototype Generation
where $K$ denotes the number of Gaussian Components, and $\mu_{m}^{k}$, $\Sigma_{m}^{k}$ represent the mean and covariance of the $k$-th Gaussian component for modality $m$, respectively. Then the probability of $n$-th sample $\mathbf{x}_{n}$ belonging to the $k$-th Gaussian component is calculated as:
@@
3.2.2 Prototype-guided Optimal Transport
The modality-unique features of different modalities often lie in distinct feature spaces with significant distributional differences, traditional point-to-point alignment methods struggle to capture both global and local relationships. To address this challenge in multimodal scenarios, we introduce a multi-marginal Optimal Transport approach to establish matches between distributions. The cross-modal prototype matching cost matrix is defined as:
@@ 3.2.2 Prototype-guided Optimal Transport
$C_{i,j}(k_{i},k_{j})=||\mu_{i}^{k_{i}}-\mu_{j}^{k_{j}}||^{2}+\text{Tr}(\Sigma_{i}^{k_{i}}+\Sigma_{j}^{k_{j}}-2(\Sigma_{i}^{k_{i}}\Sigma_{j}^{k_{j}})^{\frac{1}{2}})$ (6)
The optimization objective for cross-modal prototype alignment aims to minimize the total alignment cost across all modalities while satisfying marginal distribution constraints. The objective function is defined as:
$T^{*}=\arg\min_{T}\sum_{k}T(k)\cdot C(k)+\lambda\sum_{k}T(k)\log T(k)$ (7)
@@
$\sum_{k_{j}:j\neq i}T(k_{1},k_{2},\dots,k_{M})=\nu_{i}(k_{i}),\forall i\in\{1,2,\dots,M\},\forall k_{i}$ (8)
$\nu_{i}(j_{i})$ represents the marginal distribution of modality $m_{i}$ over its prototypes. Combining global alignment via Optimal Transport and local alignment through sample-to-prototype calibration, the overall heterogeneity alignment loss is defined as:
$\mathcal{L}_{hete}=\sum_{k}T^{*}(k)\cdot C(k)+\frac{1}{N}\sum_{n=1}^{N}\sum_{k=1}^{K}w_{i}^{n}(k)\cdot||\mathcal{F}_{i}^{n}-\mu_{j\neq i}^{k}||^{2}$ (9)
the first term, $\mathcal{L}_{OT}$ aligns the distributions of prototypes across modalities, ensuring global consistency. The second term $\mathcal{L}_{Proto}$ ensures fine-grained alignment by minimizing the weighted distance between samples $x_{i}^{n}$ in source modality $i$ and prototypes in target modality $j$. By combining $\mathcal{L}_{OT}$ and $\mathcal{L}_{Proto}$, this heterogeneous alignment loss captures both global and local relationships, providing a robust mechanism for aligning heterogeneous modalities in a unified feature space.

### v2

@@ Abstract
Multimodal representation learning aims to capture both shared and complementary semantic information across multiple modalities. This intrinsic heterogeneity of diverse modalities presents substantial challenges to achieving effective cross-modal collaboration and integration. To address this, we introduce DecAlign, a novel hierarchical cross-modal alignment framework designed to decouple multimodal representations into modality-unique (heterogeneous) and modality-common (homogeneous) features. Specifically, we mitigate distributional discrepancies for modality-unique features via a novel prototype-guided optimal transport alignment strategy leveraging Gaussian mixture modeling and multi-marginal transport. Concurrently, semantic consistency across modalities is reinforced by aligning latent distribution matching with maximum mean discrepancy regularization. Furthermore, we incorporate a multimodal transformer to enhance high-level semantic feature fusion, further reducing cross-modal inconsistencies. Our extensive experiments across four widely used multimodal benchmarks demonstrate that DecAlign consistently outperforms state-of-the-art methods on five metrics. These results highlight the efficacy of DecAlign in improving cross-modal alignment and semantic consistency while preserving modality-unique features, marking a significant advancement in multimodal representation learning scenarios. Our project page is at https://taco-group.github.io/DecAlign/.
@@ 1 Introduction
Multimodal representation learning seeks to effectively integrate diverse modalities by capturing their shared semantics while retaining modality-unique characteristics. This goal has been pursued across numerous domains, including multimodal sentiment analysis (Lian et al., 2023; Das & Singh, 2023; Wang et al., 2024a), recommendation systems (Liu et al., 2024a; 2022), autonomous driving (Yuan et al., 2025a; Xing et al., 2024b; Ma et al., 2025; Xing et al., 2024a), out-of-distribution detection (Dong et al., 2024; Li et al., 2024c), and general visual understanding and reasoning (Xing et al., 2025; Wang et al., 2024b; Cui et al., 2023; Wang et al., 2022b; Liang et al., 2025). Despite significant advancements, the intrinsic heterogeneity among modalities—mainly due to divergent data distributions, various representation scales, and semantic granularities—remains a critical barrier that hampers effective cross-modal integration.
Motivation. This challenge is further intensified by the complex entanglement of modality-unique (heterogeneous) patterns and cross-modal common (homogeneous) semantics. Conventional multimodal fusion methods typically simplify the issue by projecting raw multimodal data into unified spaces via straightforward concatenation or linear transformations(Han et al., 2022; Zhang et al., 2023). However, this indiscriminate fusion often entangles modality-unique features with global shared semantics, leading to semantic interference, wherein detailed unimodal characteristics may disrupt global cross-modal relationships (Liang et al., 2024a; Xu et al., 2023). This phenomenon is particularly evident when dealing with dimensional mismatches, such as high-dimensional, spatially correlated image features paired with low-dimensional, temporally correlated text features (Wei et al., 2025; 2024; Zhu et al., 2024). These dimensional mismatches frequently lead to suboptimal alignment, causing either information redundancy or critical loss during fusion.
@@
Figure 1: DecAlign achieves superior performance compared to state-of-the-art methods across multiple multimodal benchmarks. The bubble size represents relative model performance, illustrating the trade-off between Acc-2 and Binary F1 Score.
Our Approach. To overcome these limitations, we propose DecAlign, a hierarchical cross-modal alignment framework for multimodal representation learning. As illustrated in Figure 2, DecAlign first explicitly decouples heterogeneous and homogeneous features through specialized encoders. Then, leveraging a dual-stream cross-modal alignment mechanism, DecAlign individually handles modality characteristics at different granularities: ❶ For heterogeneity, we propose prototype-based optimal transport alignment (Peyré & Cuturi, 2019) using Gaussian Mixture Modeling (GMM) (Bishop, 2006) and multi-marginal transport plans (Pass, 2015), effectively mitigating distribution discrepancies and constraining modality-unique interference. Additionally, we enhance semantic alignment and robustness through a multimodal transformer, which employs cross-modal attention mechanisms to bridge high-level semantic inconsistencies. ❷ For homogeneity, DecAlign achieves semantic consistency via latent distribution matching with Maximum Mean Discrepancy (MMD) regularization. Finally, we concatenate the aligned modality-unique features with modality-common features, passing them through a learnable projector for downstream tasks. Our key contributions are summarized as follows:
@@
  • Empirical Evaluation. Extensive experiments on four widely used benchmark datasets demonstrate that DecAlign consistently outperforms 13 state-of-the-art methods, validating its efficacy and generalizability for multimodal representation learning.
2 Related Work (Extended Version in Appendix A)
Multimodal Representation Learning. This field integrates heterogeneous modalities into unified representations that capture complementary semantics (Qian et al., 2025b; Liang et al., 2024b; Bayoudh, 2024; Wang et al., 2025). Advances include contrastive and masked modeling (Self-MM), and hierarchical graph contrastive learning (HGraph-CL) (Yu et al., 2021; Lin et al., 2022). Yet entanglement of heterogeneity and complementarity hampers leveraging both. To address this, MISA disentangles invariant and unique features, while DMD applies graph knowledge distillation (Hazarika et al., 2020; Li et al., 2023). However, global modeling dominates, often neglecting token-level inconsistencies. Our DecAlign introduces hierarchical alignment, moving from local to global, heterogeneity to homogeneity, for precise and consistent integration.
@@ 2 Related Work (Extended Version in Appendix A)
Cross-Modal Alignment. The core challenge in multimodal learning is structural, distributional, and semantic heterogeneity, which restricts feature synergy (Zhu et al., 2024). Main approaches include: ❶ Shared Representation. Learning a unified latent space for semantic consistency. CLIP aligns image–text pairs via large-scale contrastive learning (Radford et al., 2021; Gao et al., 2024), while Uni-Code uses disentangling and exponential moving average for stable alignment (Xia et al., 2024b). ❷ Transformer-based Cross-Attention. Cross-attention dynamically captures information across modalities, as in multimodal transformers with disentangled or hierarchical fusion (Tsai et al., 2019; Yang et al., 2022; Hu et al., 2024). ❸ Modality Translation. Translation methods build mappings through cross-modal generation or reconstruction, explicitly modeling dependencies (Liu et al., 2024b; Zeng et al., 2024; Tian et al., 2022). ❹ Knowledge Distillation. Distillation balances inter-modal contributions by transferring knowledge. DMD applies graph distillation for correlation modeling, and UMDF uses unified self-distillation for robust representation learning (Li et al., 2023; 2024b). Compared with methods that risk over-alignment and loss of modality-specific traits, our framework combines representation decoupling with hierarchical alignment to preserve unimodal uniqueness while ensuring semantic consistency.
Figure 2: The framework of our proposed DecAlign approach, illustrated in a multimodal setting with visual, audio, and language inputs. Modality Feature Encoders first extract unimodal embeddings, which are then decoupled into modality heterogeneous and homogeneous components by modality-unique/common encoders. Heterogeneous features are aligned via optimal transport-based cross-modal prototypes, and homogeneous semantics are aligned through latent space semantics and Maximum Mean Discrepancy-based distribution matching. Heterogeneous features are refined by a multimodal transformer for capturing finer-grained cross-modal interactions, then concatenated with homogeneous features and passed through a fully connected layer for downstream tasks.
3 Method
@@ 3.1 Multimodal Feature Decoupling
Given a multimodal dataset with $M$ modalities, each modality $m$ provides features with its unique temporal length $T_{m}$ and feature dimension $d_{m}$. Due to this inherent variation across modalities, we apply modality-unique 1D temporal convolution layers that aggregate local temporal patterns and transform all features to the same temporal length $T_{s}$ and feature dimension $d_{s}$. The resulting unimodal features are expressed as: $\tilde{\mathbf{X}}_{m}\in\mathbb{R}^{T_{s}\times d_{s}}$. The primary challenge in multimodal tasks lies in the inherent heterogeneity across modalities, hindering the integration of homogeneous features. To address this, we decouple the multimodal representations into modality-common features, which emphasize semantic consistency across modalities, and modality-unique features, capturing modality-unique characteristics with some redundancy. Building upon this, we employ three modality-unique encoders $\mathbf{E}_{\mathrm{uni}}^{(m)}$ and a modality-shared encoder $\mathbf{E}_{\mathrm{com}}$, to extract heterogeneous features as $\mathcal{F}_{\mathrm{uni}}^{(m)}=\mathbf{E}_{\mathrm{uni}}^{(m)}(\tilde{\mathbf{X}}_{m})$ and cross-modal homogeneous features as $\mathcal{F}_{\mathrm{com}}^{(m)}=\mathbf{E}_{\mathrm{com}}(\tilde{\mathbf{X}}_{m})$.
Considering the inherent heterogeneity and potential redundancy across modalities, we refine the decoupling process by explicitly separating modality-unique and modality-common features. All encoders are designed to produce representations with the same dimensionality to ensure compatibility. Instead of modeling distributions or computing mutual information which can be computationally expensive, we use cosine similarity to quantify their potential overlap. Hence, the loss of decoupling process is formally defined as:
[…]
@@ 3.2 Heterogeneity Alignment
In multimodal tasks, modality-unique features capture distinct characteristics specific to each modality. However, these features often differ significantly in spatial structure, scale, noise level, and density, making direct point-to-point alignment across modalities both unreliable and computationally expensive. Moreover, although these features vary in form, they frequently carry semantically aligned information when referring to the same underlying concept or object category. To effectively bridge modality-unique feature differences while preserving shared semantic structures, we introduce category prototypes as semantic anchors across modalities. These prototypes represent consistent semantic patterns underlying different modality-specific representations and serve as reference points to guide alignment. Building on this, we employ a prototype-guided multi-marginal optimal transport framework to achieve adaptive and fine-grained alignment across heterogeneous feature spaces.
Prototype Generation. To flexibly capture the complex distributions and potential correlations in multimodal data, we employ the Gaussian Mixture Model (GMM), which leverages its soft assignment mechanism and Gaussian distribution assumption to more accurately represent the prototype structures of different modality features. The GMM is fitted using the standard Expectation-Maximization algorithm, which iteratively estimates the mixture coefficients, means, and covariances to maximize the likelihood of the modality-unique features. We first model modality-unique features using GMM, with prototypes represented by the mean and covariance of Gaussian distributions:
@@
where $K$ denotes the number of Gaussian components, which is set equal to the category number in downstream task, and $\mu_{m}^{k}$, $\Sigma_{m}^{k}$ represent the mean and covariance of the $k$-th Gaussian component for modality $m$, respectively. Then the probability of $n$-th sample $\mathbf{x}_{n}$ belonging to the $k$-th Gaussian component is calculated as:
@@
Prototype-guided Optimal Transport. The modality-unique features of different modalities often lie in distinct feature spaces with significant distributional differences, traditional point-to-point alignment methods struggle to capture both global and local relationships. To address this challenge in multimodal scenarios, we introduce a multi-marginal Optimal Transport approach to establish matches between distributions. The cross-modal prototype matching cost matrix is defined as:
@@
[…]
The optimization objective for cross-modal prototype alignment aims to minimize the total alignment cost across all modalities while satisfying marginal distribution constraints. The objective function is
$T^{*}=\arg\min_{T}\sum_{k}T(k)\cdot C(k)+\lambda\sum_{k}T(k)\log T(k),$ (7)
@@
$\sum_{k_{j}:j\neq i}T(k_{1},k_{2},\dots,k_{M})=\nu_{i}(k_{i}),\forall i\in\{1,2,\dots,M\},\forall k_{i},$ (8)
where $\nu_{i}(j_{i})$ represents the marginal distribution of modality $m_{i}$ over its prototypes. Combining global alignment via Optimal Transport and local alignment through sample-to-prototype calibration, the overall heterogeneity alignment loss is defined as:
[…]
The first term, $\mathcal{L}_{OT}$, aligns the distributions of prototypes across modalities, ensuring global consistency. The second term $\mathcal{L}_{Proto}$ ensures fine-grained alignment by minimizing the weighted distance between samples $x_{i}^{n}$ in source modality $i$ and prototypes in target modality $j$. By combining $\mathcal{L}_{OT}$ and $\mathcal{L}_{Proto}$, this heterogeneous alignment loss captures both global and local relationships, providing a robust mechanism for aligning heterogeneous modalities in a unified feature space.